Monday, 24 August

Monday, 24 August

Invited Lectures

\underline{14:00 – 14:45}

02 Sergei Ivanov, University of Illinois, Urbana, USA On the Burnside problem for groups of even exponent (I: p.38) Chair: Otto H. Kegel

MA 005

05 Ulrich Pinkall, TU Berlin, Berlin, Germany Quaternionic line bundles and the differential geometry of surfaces (I: p.75) Chair: Dirk Ferus

MA 001

10 Dmitri Yafaev, Université Rennes-1, Rennes, France Scattering theory: some old and new problems (I: p.127) Chair: Volker Enß

H 1012

12 Friedrich Götze, Universität Bielefeld, Bielefeld, Germany Lattice Point Problems and the Central Limit Theorem in Euclidean Spaces (I: p.143) Chair: Richard M. Dudley

H 1028

13 András Frank, Eötvös Lorand University, Budapest, Hungary Applications of Relaxed Submodularity (I: p.155) Chair: Alexander Schrijver

H 2032

19 Jeremy John Gray, The Open University, Milton Keynes, England The Riemann-Roch Theorem and Geometry, 1854-1914 (I: p.208) Chair: Umberto Bottazzini

MA 004

\underline{15:00 – 15:45}

02 Idun Reiten, Norwegian University of Science and Technology, Trondheim, Norway Tilting Theory and Quasitilted algebras (I: p.43) Chair: Claus Michael Ringel

MA 005

05 Tobias Holck Colding, Courant Institute, New York University, New York, USA Manifolds and spaces with Ricci curvature bounds (I: p.68) Chair: Dirk Ferus

MA 001

10 Hans Lindblad, University of California at San Diego, La Jolla, USA Minimal regularity solutions of nonlinear wave equations (I: p.122) Chair: Konrad Gröger H 1012 12 & 16 Marco Avellaneda, Courant Institute, New York University, New York, USA Inverse Problems in Quantitative Financial Modeling (I: p.140) Chair: George C. Papanicolaou

H 1028

13 Jiří Matoušek, Charles University, Praha, Czech Republic Mathematical snapshots from the computational geometry landscape (I: p.158) Chair: Jaroslav Nešetril

H 2032

19 Karine Chemla, CNRS and Université Paris 7, Paris, France History of Mathematics in China: A factor in World History and a Source for New Questions (I: p.206) Chair: Sergei S. Demidov

MA 004

\underline{16:15 – 17:00}

03 François Gramain, Université Jean Monnet Saint-Etienne, Saint-Etienne, France Some results of algebraic independence (I: p.51) Chair: Michel Waldschmidt

MA 005

08 Stephen Wainger, University of Wisconsin, Madison, USA Discrete Analogues of Singular and Maximal Radon Transforms (I: p.109) Chair: Detlef H. Müller

MA 001

09 Grzegorz Świa̧tek, Pennsylvania State University, University Park, USA Induced Hyperbolicity for One-Dimensional Maps (I: p.119) Chair: Klaus R. Schneider

H 1012

15 Hisashi Okamoto, RIMS, Kyoto University, Kyoto, Japan Bifurcation of Kolmogorov flows with an emphasis on singular limit (I: p.180) Chair: Rolf Jeltsch

H 1028

16 Gérard M. Iooss, Institut Universitaire de France, Valbonne, France Travelling water-waves, as a paradigm for bifurcations in reversible infinite dimensional “dynamical” systems (I: p.184) Chair: Klaus Kirchgässner

H 2032

19 Joseph Dauben, The City University of New York, New York, USA Marx, Mao and Mathematics: the Story of Nonstandard Analysis in China and the Cultural Revolution (I: p.208) Chair: Ivor Grattan-Guiness

MA 004

\underline{17:15 – 18:00}

03 William Duke, Rutgers University, New Brunswick, USA Bounds for arithmetic multiplicities (I: p.50) Chair: Per Salberger

MA 005

08 Vitali Milman, Tel Aviv University, Tel Aviv, Israel Randomness and Pattern in Convex Geometric Analysis (I: p.101) Chair: Hermann König

MA 001

09 Sergei B. Kuksin, Steklov Institute, Moscow, Russia and Heriot-Watt University, Edinburgh, UK Towards a Qualitative Theory of Hamiltonian PDEs (I: p.116) Chair: Jürgen Moser

H 1012

15 Gregory Beylkin, University of Colorado at Boulder, Boulder, USA On multiresolution methods in numerical analysis (I: p.178) Chair: Siegfried Prößdorf

H 1028

16 Ulf Grenander, Brown University, Providence, USA Strategies for Seeing (I: p.183) Chair: Friedrich Götze

H 2032

18 Mogens Niss, Roskilde University, Roskilde, Denmark Aspects of the nature and the state of research in mathematical education (I: p.198) Chair: Urs Kirchgraber MA 004